# Strong coupling non-Markovian quantum thermodynamics of a finite-bath system

### v2

@@ V Ergotropy: application of the central spin model as a quantum battery
$\rho_{S}^{p}=\sum_{j}r_{j}\ket{\epsilon_{j}}\bra{\epsilon_{j}},$
$H_{S}=\sum_{k}\epsilon_{k}\ket{\epsilon_{k}}\bra{\epsilon_{k}},$ (64)
where the eigenvalues $r_{j}$ and $\epsilon_{j}$ have the following ordering: $r_{1}\geq r_{2}\geq r_{3}...\text{ as }\epsilon_{1}\leq\epsilon_{2}\leq\epsilon_{3}...$ Thus, the passive state is defined in such a way that the minimum energy state has the maximum population. For the central spin system (or, for any single qubit system with $H_{S}=\frac{\omega_{0}}{2}\sigma_{z}$ [37]), using the Bloch vector form of the system’s state at any time $t$, discussed in Eq. (39), the expression for the ergotropy is given by
@@
The ergotropy can be further divided into its incoherent ($\mathcal{W}_{i}$) and coherent parts ($\mathcal{W}_{C}$[101], such that $\mathcal{W}=\mathcal{W}_{i}+\mathcal{W}_{C}$. The incoherent ergotropy can be calculated in a similar manner as the ergotropy of the system by replacing the system’s state $\rho_{S}(t)$ in Eq. (63) with a dephased state $\rho_{S}^{D}(t)$. From Eq. (39), the dephased state is given by $\rho_{S}^{D}(t)=\frac{1}{2}\left\{[1+z(t)]\ket{0}\bra{0}+[1-z(t)]\ket{1}\bra{1}\right\}$ | }, where |$\ket{0}(\ket{1})$ denotes the excited (ground) state of the system. The passive state $\rho_{S}^{D,p}$ corresponding to this dephased state depends on the sign of $z(t)$. For the single qubit systems with Hamiltonian $H_{S}=\frac{\omega_{0}}{2}\sigma_{z}$ and with positive $z(t)$, the incoherent ergotropy boils down to $\mathcal{W}_{i}[\rho_{S}(t)]=\mathcal{W}[\rho_{S}^{D}(t)]=\omega_{0}z(t)$, which for negative $z(t)$ becomes zero. Thus, the coherent ergotropy $\mathcal{W}_{C}[\rho_{S}(t)]$ is equal to the ergotropy $\mathcal{W}[\rho_{S}(t)]$ for negative $z(t)$ and $\mathcal{W}_{C}[\rho_{S}(t)]=\frac{\omega_{0}}{2}\left[\sqrt{x(t)^{2}+y(t)^{2}+z(t)^{2}}-z(t)\right]$ for $z(t)$ being positive.
@@
Ergotropy serves as an important quantifier for the charging and discharging behavior of the system. The system is said to be charging (discharging) when its ergotropy increases (decreases). To this end, the central spin can be visualized as a quantum battery, with the spin bath acting as a charger. A similar analysis in the case of open quantum systems with the system being surrounded by a bosonic bath has been done in [37, 38], where the system was initially taken to be in a positive ergotropy state and the bath acted as a re-charger (due to non-Markovian evolution). However, in the case of the spin bath considered here, we seek whether the bath can act as a charger, that is, whether the bath itself can lead to a state giving some positive ergotropy, even if we take an initial state with zero ergotropy. To check this, let us consider the central spin system to be initially in a ground state $\rho_{S}(0)=\ket{1}\bra{1}$ |, which is a passive state, from Eq. (63). For this initial state, the factors $x(t)$ and $y(t)$ from Eq. (39) remain zero at any time $t$. The ergotropy $\mathcal{W}[\rho_{S}(t)]$ in this case is given by
$\mathcal{W}[\rho_{S}(t)]=\frac{\omega_{0}}{2}\left[z(t)+\left|z(t)\right|\right]~{}~{}~{}\text{for}~{}~{}\rho_{S}(0)=\ket{1}\bra{1}.$ | . (66)
@@
In the case of the ground state, the value of the $z(0)=-1$, and therefore, $\mathcal{W}[\rho_{S}(0)]=0$. However, the factor $z(t)$ can become positive, thereby making ergotropy positive as time progresses. The condition on $z(t)>0$ leads to $\eta(t)>\frac{1}{2}$, with the help of Eq. (IV.1).
@@
Figure 8 shows the variation of the $\eta(t)$ with respect to time. It is observed that $\eta(t)$ crosses the $1/2$ line when the interaction strength $\epsilon$ is increased. Further, it is interesting to note that upon increasing the temperature of the bath, $\eta(t)$ is greater than $1/2$ at multiple times. This suggests that the tendency of the bath to act as a charger increases as temperature increases. The number of spins in the bath $N$ also favors the role of the bath as a charger, as in this case, too, the number of times $\eta(t)$ is greater than $1/2$ increases with an increase in $N$, as can be seen from Fig. 8(c). Figure 8(d) is an example where the initial state of the system has zero ergotropy, yet the ergotropy revives multiple times throughout the evolution due to the bath acting as a charger.
@@ Appendix A The spectral decomposition of the total Hamiltonian
|$\ket{i}$ is the standard computational basis in the above equations. The factor $\chi_{\pm}(n)$ is given by

### v3

@@ V Ergotropy: application of the central spin model as a quantum battery
$|,\displaystyle\rho_{S}^{p}=\sum_{j}r_{j}\ket{\epsilon_{j}}\bra{\epsilon_{j}}$,
$|,\displaystyle H_{S}=\sum_{k}\epsilon_{k}\ket{\epsilon_{k}}\bra{\epsilon_{k}}$, (64)
where the eigenvalues $r_{j}$ and $\epsilon_{j}$ have the following ordering: $r_{1}\geq r_{2}\geq r_{3}...\text{ as }\epsilon_{1}\leq\epsilon_{2}\leq\epsilon_{3}$... . Thus, the passive state is defined in such a way that the minimum energy state has the maximum population. For the central spin system (or, for any single qubit system with $H_{S}=\frac{\omega_{0}}{2}\sigma_{z}$ [37]), using the Bloch vector form of the system’s state at any time $t$, discussed in Eq. (39), the expression for the ergotropy is given by
@@
[…]
@@
Ergotropy serves as an important quantifier for the charging and discharging behavior of the system. The system is said to be charging (discharging) when its ergotropy increases (decreases). To this end, the central spin can be visualized as a quantum battery, with the spin bath acting as a charger. A similar analysis in the case of open quantum systems with the system being surrounded by a bosonic bath has been done in [37, 38], where the system was initially taken to be in a positive ergotropy state and the bath acted as a re-charger (due to non-Markovian evolution). However, in the case of the spin bath considered here, we seek whether the bath can act as a charger, that is, whether the bath itself can lead to a state giving some positive ergotropy, even if we take an initial state with zero ergotropy. To check this, let us consider the central spin system to be initially in a ground state $|\rho_{S}(0)=\ket{1}\bra{1}$, which is a passive state, from Eq. (63). For this initial state, the factors $x(t)$ and $y(t)$ from Eq. (39) remain zero at any time $t$. The ergotropy $\mathcal{W}[\rho_{S}(t)]$ in this case is given by
$|.\displaystyle\mathcal{W}[\rho_{S}(t)]=\frac{\omega_{0}}{2}\left[z(t)+\left|z(t)\right|\right]~~~\text{for}~~\rho_{S}(0)=\ket{1}\bra{1}$. (66)
@@
In the case of the ground state, the value of z(0)=1z(0)=-1, and therefore, $\mathcal{W}[\rho_{S}(0)]=0$. However, the factor $z(t)$ can become positive, thereby making ergotropy positive as time progresses. The condition on $z(t)>0$ leads to $\eta(t)>\frac{1}{2}$, with the help of Eq. (IV.1).
@@
Figure 8 shows the variation of the $\eta(t)$ with respect to time. It is observed that $\eta(t)$ crosses the $1/2$ line when the interaction strength $\epsilon$ is increased. Further, it is interesting to note that upon increasing the temperature of the bath, $\eta(t)$ is greater than $1/2$ at multiple times. This suggests that the tendency of the bath to act as a charger increases as the temperature increases. The number of spins in the bath $N$ also favors the role of the bath as a charger, as in this case, too, the number of times $\eta(t)$ is greater than $1/2$ increases with an increase in $N$, as can be seen from Fig. 8(c). Figure 8(d) is an example where the initial state of the system has zero ergotropy, yet the ergotropy revives multiple times throughout the evolution due to the bath acting as a charger.
@@ Appendix A The spectral decomposition of the total Hamiltonian
[…]